### v4

# A Systematic Review of Aspect-based Sentiment Analysis (ABSA): Domains, Methods, and Trends thanks: Preprint manuscript

Abstract
Aspect-based Sentiment Analysis (ABSA) is a fine-grained type of sentiment analysis that identifies aspects and their associated opinions from a given text. With the surge of digital opinionated text data, ABSA gained increasing popularity for its ability to mine more detailed and targeted insights. Many review papers on ABSA subtasks and solution methodologies exist, however, few focus on trends over time or systemic issues relating to research application domains, datasets, and solution approaches. To fill the gap, this paper presents a Systematic Literature Review (SLR) of ABSA studies with a focus on trends and high-level relationships among these fundamental components. This review is one of the largest SLRs on ABSA, and also, to our knowledge, the first that systematically examines the trends and inter-relations among ABSA research and data distribution across domains and solution paradigms and approaches. Our sample includes 519 primary studies screened from 4191 search results without time constraints via an innovative automatic filtering process. Our quantitative analysis not only identifies trends in nearly two decades of ABSA research development but also unveils a systemic lack of dataset and domain diversity as well as domain mismatch that may hinder the development of future ABSA research. We discuss these findings and their implications and propose suggestions for future research.
Keywords ABSA, aspect-based sentiment analysis, systematic literature review, natural language processing, domain
1 Introduction
In the digital era, a vast amount of online opinionated text is generated daily through which people express views and feelings (i.e. sentiment) towards certain subjects, such as user reviews, social media posts, and open-ended survey question responses [1]. Understanding the sentiment of these opinionated text data is essential for gaining insights into people’s preferences and behaviours and supporting decision-making across a wide variety of domains [2, 3, 4, 5, 6]. The analyses of opinionated text usually aim at answering questions such as “What subjects were mentioned?”, “What did people think of (a specific subject)?”, and “How are the subjects and/or opinions distributed across the sample?” (e.g. [7, 8, 9, 10]). These objectives, along with today’s enormous volume of digital opinionated text, require an automated solution for identifying, extracting and classifying the subjects and their associated opinions from the raw text.
Aspect-based sentiment analysis (ABSA) is one such solution. ABSA is a sub-field of Sentiment Analysis (SA), which is a core task of natural language processing (NLP). SA, also known as “opinion mining” [5, 6, 11, 12, 13], solves the problem of identifying and classifying given text corpora’s affect or sentiment orientation [14, 4] into polarity categories (e.g. “positive, neutral, negative”) [15, 16, 17], intensity/strength scores (e.g. from 1 to 5) [18], or other categories. The “identifying the subjects of opinions” part of the quest relates to the granularity of SA. Traditional SA mostly focuses on document- or sentence-level sentiment and thus assumes a single subject of opinions [19, 20]. In recent decades, the explosion of online opinion text has attracted increasing interest in distilling more targeted insights on specific entities or their aspects within each sentence through finer-grained SA [19, 20, 14, 21, 22]. This is the problem Aspect-based sentiment analysis (ABSA) aims to solve.
A complete ABSA solution that identifies aspects and their corresponding sentiments typically consists of a combination of subtasks detailed in Subsection 2.1.1. Subtask combinations can, to a large extent, reflect different ABSA problem formulations and tactics. The techniques of solving these subtasks evolved from linguistic and statistical to the dominant machine learning (ML) approaches [23, 24, 20, 25], with Deep Learning (DL) methods on the rise in recent years [20, 26].
1.1 Review Rationale
This review is underpinned by the following rationales:
First, the shift towards ML, especially supervised and/or DL solutions for ABSA, highlights the importance of dataset resources, particularly annotated, large benchmark datasets for the quality and development of ABSA research. Meanwhile, the finer granularity of ABSA also brings the persistent challenge of domain dependency described in Subsection 2.1.2. The diversity of ABSA datasets and their domains can have a direct and systematic impact on research and applications.
The early seminal works in ABSA were motivated by commercial applications and focused on product and service reviews [20, 27, 26], such as Ganu et al., (2009) [28], Hu and Liu (2004a [29], 2004b [30]), and Pontiki et al. (2014 [31], 2015 [32], 2016 [33]) that laid influential foundations with widely-used product and service review ABSA benchmark datasets [27, 26]. However, the need for mining insights from opinions far exceeds this single domain. Many other areas, especially the public sector, also have an abundance of opinionated text data and can benefit from ABSA, such as helping policy-makers understand public attitudes and reactions towards events or changes [2], improving healthcare services and treatments via patient experience and concerns in clinical visits, symptoms, drug efficacy and side-effects [34, 35], and guiding educators in meeting teacher and learner needs and improving their experience [3, 4, 5, 6]. While the more general SA research has been applied to “nearly every domain” ( [19], P.1), this does not seem to be the case for ABSA. Chebolu et al. (2022) reviewed 62 public ABSA datasets released between 2004 and 2020 covering “over 25 domains” ( [36], p.1). However, 53 out of these 62 datasets were reviews of restaurants, hotels, and digital products; only five were not related to commercial products or services, and merely one was on the public sector domain (university reviews). Will this dataset domain homogeneity be found with a larger sample of primary studies? Does this domain skewness reflect the concentration of ABSA research focus or merely the lack of dataset diversity? Answers to these questions could inform and shape future ABSA research through individual research decisions and community resource collaboration.
Second, there are many good survey papers on ABSA, most focused on introducing the methodological details of common ABSA subtasks and solutions (e.g. [23, 37, 27, 38, 39, 40] or specific approaches such as DL methods for ABSA (e.g. [20, 26, 41, 42]). However, many of these reviews only explored each subtask and/or technique individually, and few examined their combinations or changes over time. For example, although the above-listed reviews [20, 26, 41, 42] reported the rise of DL approaches in ABSA similar to that of NLP as a whole, it is unclear whether ABSA research was also increasingly dominated by the attention mechanism from the Transformer architecture [43] and pre-trained large language models since 2018 [44], and if linguistic and traditional ML approaches were still active. In addition, most of these surveys used a smaller and selected sample that could not support conclusions on trends. As the field matures, we believe it is necessary and important to examine trends and matters outside the problem solution itself, so as to inform research decisions, identify issues, and call for necessary community awareness and actions.
1.2 Review Focus and Research Questions
This work presents a Systematic Literature Review (SLR) of existing ABSA studies with a large-scale sample and quantitative results. We focus on trends and high-level patterns instead of methodological details that were well covered by the existing surveys mentioned above. We aim to benefit both ABSA newcomers by introducing the basics of the topic, as well as existing ABSA researchers by sharing perspectives and findings that are useful to the ABSA community and can only be obtained beyond the immediate research tasks and technicalities.
We seek to answer the following sets of Research Questions (RQs):
  1. RQ1. To what extent is ABSA research and its dataset resources dominated by the commercial (especially the product and service review) domain? What proportion of ABSA research focuses on other domains and dataset resources?
  2. RQ2. What are the most common ABSA problem formulations via subtask combinations, and what proportion of ABSA studies only focus on a specific sub-task?
  3. RQ3. What is the trend in the ABSA solution approaches over time? Are linguistic and traditional machine-learning approaches still in use?
In order to identify patterns and trends for our RQs, a sufficiently sized representative sample and systematic approach are required. We chose to conduct an SLR, as this type of review aims to answer specific research questions from all available primary research evidence following well-defined review protocols [45]. Moreover, none of the existing SLRs on ABSA share the same focus and RQs as ours: Among the 192 survey/review papers obtained from four major digital database searches detailed in Section 3, only eight were SLRs on ABSA, within which four focused on non-English language(s) [46, 47, 48, 49], two on specific domains (software development, social media) [24, 50], one on a single subtask [23], and one mentioned ABSA subtasks as a side-note under the main topic of SA [51].
This review makes a number of unique contributions to the ABSA research field: 1) It is one of the largest scoped SLRs on ABSA, with 519 eligible studies filtered from 4191 search results without time constraints; 2) To our knowledge, it is the first SLR that systematically examines the ABSA data resource distribution in relation to research application domains and methodologies; and 3) Our review methodology adopted an innovative automatic filtering process based on PDF-mining, which enhanced screening quality and reliability. Our quantitative results not only revealed trends in nearly two decades of ABSA research literature but also highlighted potential systemic issues that could limit the development of future ABSA research.
1.3 Organisation of This Review
In the following text, we introduce ABSA including its subtasks and main challenges in Section 2, describe our SLR procedures in Section 3 (“Methods”), and answer the research questions using the SLR results in Section 4 (“Results”). Finally, we discuss the key findings and acknowledge limitations in Sections 5 and 6 (“Discussion” and “Conclusion”). Appendix A contains full details of the Methods section, and Appendix B provides additional figures from the Results section.
2 Aspect-Based Sentiment Analysis (ABSA)
Aspect-Based Sentiment Analysis (ABSA) is a sub-domain of fine-grained SA [52]. ABSA focuses on identifying the sentiments towards specific entities or their attributes/ features called aspects [52, 14]. An aspect can be explicitly expressed in the text (explicit aspect) or absent from the text but implied from the context (implicit aspects) [23, 53]. Moreover, the aspect-level sentiment could differ across aspects and be different from the overall sentiment of the sentence or the document (e.g. [14, 54, 55]). Some studies further distinguish aspect into aspect term and aspect category, with the former referring to the aspect expression in the input text (e.g. “pizza”), and the latter a latent construct that is usually a high-level category across aspect terms (e.g. “food”) that are either identified or given [56, 57].
The following examples illustrate the ABSA terminologies:
Example 1 (from a restaurant review [^1] https://alt.qcri.org/semeval2014/task4/): “The restaurant was expensive, but the menu was great.
This sentence has one explicit aspect “menu” (sentiment term: “great”, sentiment polarity: positive), one implicit aspect “price” (sentiment term: “expensive”, sentiment polarity: negative). Depending on the target/given categories, the aspects can be further classified into categories, such as “menu” into “general” and “price” into “price”.
Example 2 (from a laptop review [^2] https://alt.qcri.org/semeval2015/task12/): “It is extremely portable and easily connects to WIFI at the library and elsewhere.
This sentence has two implicit aspects: “portability” (sentiment term: “portable”, sentiment polarity: positive), “connectivity” (sentiment term: “easily”, sentiment polarity: positive). The aspects can be further classified into categories, such as both under “laptop” (as opposed to “software” or “support”).
Example 3 (text from a course review): “It was too difficult and had an insane amount of work, I wouldn’t recommend it to new students even though the tutorial and the lecturer were really helpful.”
The two explicit aspects in Example 3 are “tutorial” and “lecturer” (sentiment terms: “helpful”, polarities: positive). The implicit aspects are “content” (sentiment term: “too difficult”, sentiment polarity: negative), “workload” (sentiment term: “insane amount”, sentiment polarity: negative), and “course” (sentiment term: “would not recommend”, sentiment polarity: negative). An illustration of aspect categories would be assigning the aspect “lecturer” to the more general category “staff” and “tutorial” to the category “course component”.
As demonstrated above, the fine granularity makes ABSA more targetable and informative than document- or sentence-level SA. Thus, ABSA can precede downstream applications such as attribute weighting in overall review ratings (e.g. [58]), aspect-based opinion summarization (e.g. [59, 60, 61], and automated personalised recommendation systems (e.g. [62, 63]).
2.1 ABSA Subtasks and Challenges
Compared with document- or sentence-level SA, while being the most detailed and informative, ABSA is also the most complex and challenging [64]. The most noticeable challenges include the number of ABSA subtasks, their interrelations and context dependencies, and the generalisability of solutions across topic domains.
2.1.1 ABSA Subtasks
A full ABSA solution has more subtasks than coarser-grained SA. The most fundamental ones [55, 64, 65, 66] include:
{quoting}
[vskip=12pt]
Aspect (term) Extraction/Identification (AE), which has a slight variation in meaning depending on the overall ABSA approach. Some authors (e.g. [67, 68, 69]) consider AE as identifying the attribute or entity that is the target of an opinion expressed in the text and sometimes call it “opinion target extraction” [70]. In these cases, opinion terms were often identified in order to find their target aspect terms. Others (e.g. [14, 71, 65, 22, 13]) define AE as identifying the key or all attributes of entities mentioned in the text. Implicit-Aspect Extraction (IAE) is often mentioned as a task by itself due to its technical challenge.
Opinion (term) Extraction/Identification (OE), which relates to identifying the “opinion terms” or the sentiment expression of a specific entity/aspect (e.g. [55, 72, 73, 66]). In Example 1 above, an OE task would extract the sentiment terms “great” (associated with the aspect term “menu”) and “expensive” (associated with the implicit aspect “price”).
Aspect-Sentiment Classification (ASC), which refers to obtaining the sentiment polarity category (e.g. negative, neutral, positive, conflict) or sentiment score (e.g. 1 to 5 or -1 to 1 along the scale from negative to positive) associated with a given aspect or aspect category (e.g. [14, 74, 75]). This is often done via evaluating the associated opinion term(s), and sentiment lexicon resources such as the SentiWordNet [76] and SenticNet [77] can be used to assign polarity scores [74]. Sentiment scores can be further aggregated across opinion terms for the same aspect, or across aspect terms to generate higher-level ratings, such as aspect-category ratings within or across documents [74, 75].
As an extension of AE, some studies also involve Aspect-Category Detection (ACD) and Aspect Category Sentiment Analysis (ACSA) when the focus of sentiment analysis is on (often pre-defined) latent topics or concepts and requires classifying aspect terms into categories [78].
Traditional full ABSA solutions often perform the subtasks in a pipeline manner [79, 80] using one or more of the linguistic (e.g. lexicons, syntactic rules, dependency relations), statistical (e.g. n-gram, Hidden Markov Model (HMM)), and machine-learning approaches [23, 24, 25]. For instance, for AE and OE, some studies used linguistic rules and sentiment lexicons to first identify opinion terms and then the associated aspect terms of each opinion term, or vice versa (e.g. [21, 34]), and then moved on to ASC or ACD using a supervised model or unsupervised clustering and/or ontology [63, 74]. Hybrid approaches are common given the task combinations in a pipeline.
With the rise of multi-task learning and deep learning [81], an increasing number of studies explore ABSA under an End-to-End (E2E) framework that performs multiple fundamental ABSA subtasks in one model, and some combine them into a single composite task [64, 79, 82]. These composite tasks are most commonly formulated as a sequence- or span-based tagging problem [64, 79, 80]. The most common composite tasks are: Aspect-Opinion Pair Extraction (AOPE), which directly outputs {aspect, opinion} pairs from text input [80, 83, 84] such as “⟨menu, great⟩” from Example 1; Aspect-Polarity Co-Extraction (APCE) [64, 85], which outputs {aspect, sentiment polarity} pairs such as “⟨menu, positive⟩”; Aspect-Sentiment Triplet Extraction (ASTE) [64, 79, 86, 66], which outputs {aspect, opinion, sentiment category} triplets, such as “⟨menu, great, positive⟩”; and Aspect-Sentiment Quadraplet Extraction (ASQE) [87] that outputs {aspect, opinion, aspect category, sentiment category} quadruplets, such as “⟨menu, great, general, positive⟩”.
As this review focuses on trends instead of detailed solutions and methodologies, we refer interested readers to existing review papers that provide a good summary of common ABSA subtask solutions and approaches, such as AE (e.g. [23, 27, 38, 39]), deep learning (DL) approaches for ASC (e.g. [40]), and DL methods for ABSA (e.g. [20, 26, 41, 42]).
2.1.2 Context and Domain Dependencies
The nature and the interconnection of these subtasks determine that ABSA is heavily domain- and context-dependent [19, 36, 88]. Domain refers to the ABSA task (training or application) topic domains, and context can be either the “global” context of the document or the “local” context from the text surrounding a target word token or word chunks. At least in English, the same word or phrase could mean different things or bear different sentiments depending on the context and often also the topic domains. For example, “a big fan” could be an electric appliance or a person, depending on the sentence and the domain; “cold” could be positive for ice cream but negative for customer service; and “DPS (damage per second)” could be either a gaming aspect or non-aspect in other domains. Thus, the ability to incorporate relevant context is essential for ABSA solutions; and those with zero or very small context windows, such as n-gram and Markov models, are rare in ABSA literature and can only tackle a limited range of subtasks (e.g. [89]). Moreover, although many language models (e.g. Bidirectional Encoder Representations from Transformers (BERT, [90]), Generative pre-trained transformers (GPT, [91]), Recurrent Neural Network (RNN)-based models) already incorporated local context from the input-sequence and/or general context through pre-trained embeddings, they still performed unsatisfactorily on some ABSA domains and subtasks, especially IAE, AE with multi-word aspects, AE and ACD on mixed-domain corpora, and context-dependent ASC [92, 21, 11, 88]. Many studies showed that ABSA task performance benefits from expanding the feature space beyond generic and input textual context, with domain-specific dataset/representations, and additional input features such as Part-of-Speech (POS) tags, syntactic dependency relations, lexical databases, and domain knowledge graphs or ontologies [88, 21, 11]. Nonetheless, annotated datasets and domain-specific resources are costly to produce and limited in availability, and domain adaptation, as one solution to this, has been an ongoing challenge for ABSA [93, 82, 19, 88].
The above highlights the critical role of domain-specific datasets and resources in ABSA solution quality, especially for supervised approaches. On the other hand, it suggests the possibility that the prevalence of dataset-reliant solutions in the field and a skewed ABSA dataset domain distribution could systemically hinder ABSA solution performance and generalisability [93, 94], and thus confining ABSA research and solutions close to the resource-rich domains and languages. This idea underpins this literature review’s motivation and research questions.
3 Methods
Following the guidance of Kitchenham and Charters (2007) [45], we conducted this SLR with pre-planned scope, criteria, and procedures highlighted below. The complete SLR methods and process are detailed in Appendix A.
3.1 Main Procedures
The reviewed primary studies were from four major peer-reviewed digital databases: ACM Digital Library, IEEE Xplore, Science Direct, and Springer Link. First, we manually searched and extracted 4191 database search results without publication-year constraints. Appendix Section A.1 provides more details of the search strategies and results. Next, we applied the inclusion and exclusion criteria listed in Table 1 via automatic and manual screening steps and identified 519 in-scope peer-reviewed research publications for the review. The complete screening process is summarised in Table A.2 in Section A.2. We then manually reviewed the in-scope primary studies and recorded data following a planned scheme. Lastly, we checked, cleaned, and processed the extracted data and performed quantitative analysis against our RQs.
Table 1: Inclusion and exclusion criteria used for this Systematic Literature Review (SLR)
Inclusion Criteria Exclusion Criteria
1. Published in the English language 2. Has both the sentiment analysis component and entity/aspect-level granularity 3. Focuses on text data 4. Is a primary study with quantitative elements in the ABSA approach 5. Contains original solutions or ideas for ABSA task(s) 6. Involves experiment and results on ABSA task(s) 1. The main text of the article is not in the English language 2. Missing either the sentiment analysis or entity/aspect-level granularity in the research focus 3. Only contains search keyword in the reference section 4. Only contains less than 5 ABSA-related search keywords outside the reference section 5. Is not a primary study (e.g. review articles, meta-analysis) 6. Does not provide quantitative experiment results on ABSA task(s) 7. Has fewer than three pages 8. Contains multimodal (i.e. non-text) input data for ABSA task(s) 9. The research focus is not on ABSA task(s), even though ABSA models might be involved (e.g. recommender system) 10. Focuses on transferring existing ABSA approaches between languages 11. The ABSA tasks are integrated into a model built for other purposes, and there are no stand-alone ABSA method details and/or evaluation results
3.2 Sample Summary
Figure 1 shows the number of total reviewed vs. included studies across all publication years for the 4191 search results. The search results include studies published between 1995 and 2023 (N=1), although all of the pre-2008 ones (2 from the 90s, 8 from 2003-2006, 17 from 2007) were not ABSA-focused and were excluded during automatic screening. The earliest in-scope ABSA study in the sample was published in 2008, followed by a very sparse period until 2013. The numbers of extracted and in-scope publications have both grown noticeably since 2014, a likely result of the emergence of deep learning approaches, especially sequence models such as RNNs [44, 95]. We also provide a breakdown of the included studies by publication year and type in Figure B.1 in Appendix B.
Figure 1: Number of Studies by Publication Year: Total reviewed vs. Included (N=4191)
3.3 Note on “domain” mapping
In order to answer RQ1, we made the distinction between “research application domain” (“research domain” in short) and “dataset domain”, and manually examined and classified each study and its datasets into domain categories.
We considered each study’s research domain to be “non-specific” unless the study mentioned a specific application domain or use case as its motivation. For the dataset domain, we examined each dataset used by our sample, standardised its name, and recorded the domain from which it was drawn/selected based on the description provided by the author or the dataset source webpage. Datasets without a specific domain (e.g. Twitter tweets crawled without a specific domain filter) were labelled as “non-specific”.
We then manually grouped the research and dataset domains into 19 common categories used for analysis. More details and examples on domain mapping are available in Subsection A.3 in Appendix A.
4 Results
This section presents the SLR results corresponding to each of the RQs:
4.1 Results for RQ1
RQ1. To what extent is ABSA research and its dataset resources dominated by the commercial (especially the product and service review) domain? What proportion of ABSA research focuses on other domains and dataset resources?
To answer RQ1, we examined the distribution of reviewed studies by their research (application) domains, dataset domains, and the relationship between the two. From the 519 reviewed studies, we recorded 218 datasets, 19 domain categories (15 research domains and 17 dataset domains), and obtained 1179 distinct “study - dataset” pairs and 630 unique “study - dataset domain” combinations. The key results are summarised below and presented in Table 2 and Figure 2.
Table 2: Number of In-Scope Studies Per Each Research (Application) and Dataset Domain Category
Domain Count of Studies per Research Domain % of Studies per Research Domain Count of Studies per Dataset Domain % of Studies per Dataset Domain
Non-specific 339* 65.32% 98 15.56%
Product/service review 126 24.28% 447* 70.95%
Student feedback / education review 12 2.31% 19 3.02%
Politics/ policy-reaction 8 1.54% 5 0.79%
Healthcare/ medicine 7 1.35% 9 1.43%
Video/movie review 6 1.16% 19 3.02%
News 5 0.96% 8 1.27%
Finance 5 0.96% 4 0.63%
Research/academic reviews 3 0.58% 3 0.48%
Disease 3 0.58% 3 0.48%
Employer review 1 0.19% 1 0.16%
Nuclear energy 1 0.19% 0 0.00%
Natural disaster 1 0.19% 1 0.16%
Music review 1 0.19% 1 0.16%
Multiple domain 1 0.19% 0 0.00%
Location review 0 0.00% 8 1.27%
Book review 0 0.00% 2 0.32%
Biology 0 0.00% 1 0.16%
Singer review 0 0.00% 1 0.16%
TOTAL 519 100.00% 630 100.00%
Figure 2: Distribution of Unique “Study-Dataset” Pairs (N=1179, with 519 studies and 218 datasets) by Research (Application) Domains (left) and Dataset Domains (right). Note: 1) The top flow visualises a mismatch between the two domains: the majority of studies without a specific research domain used datasets from the product/service review domain. 2) The disproportionately small number of samples in both domains that were neither “non-specific” nor “product/service review”.
In summary, our results answer RQ1 by showing that: 1) The majority (65.32%) of the reviewed studies were not for any specific application domain and only 24.28% targeted “product/service review”. 2) The dataset resources used in the sample were mostly domain-specific (84.44%) and dominated by the “product/service review” datasets (70.95%). 3) Both the research effort and dataset resources were scant in the non-commercial domains, especially the main public sector areas, with fewer than 13 studies across 14 years in each of the healthcare, policy, and education domains, where about half of the used datasets were created from scratch for the study.
Beyond RQ1, 1) and 2) above also suggest a significant mismatch between the research and dataset domains as visualised in Figure 2. Further, when filtering out datasets used by less than 10 studies, we discovered an alarming lack of dataset diversity as only 12 datasets remained, of which 10 were product/service reviews. When examining the three-way relationship among research domain, dataset domain, and dataset name, we further identified an over-representation (78.20%) of the four SemEval restaurant and laptop review benchmark datasets. This is illustrated in Figure 4.
4.1.1 Detailed Results for RQ1
For research (application) domains indicated by the stated research use case or motivation, the majority (65.32%, N=339) of the 519 reviewed studies have a “non-specific” research domain, followed by just a quarter (24.28%, N=126) in the “product/service review” category. However, the number of studies in the rest of the research domains is magnitudes smaller in comparison, with only 12 studies (2.31%) in the third largest category “student feedback/education review” since 2008, followed by 8 in Politics/policy-reaction (1.54%), and only 7 in Healthcare/medicine (1.35%). Figure 3 revealed further insights from the trend of research domain categories with five or more reviewed studies. Interestingly, “product/service review” has been a persistently major category over time, and has only been consistently taken over by “non-specific” since 2015. The sharp increase of domain-“non-specific” studies since 2018 could be partly driven by the rise of pre-trained language models such as BERT and the greater sequence processing power from the Transformer architecture and the attention mechanism [44], as more researchers explore the technicalities of ABSA solutions.
Figure 3: Number of In-Scope Studies by Research (Application) Domain and Publication Year (N=518). This graph excludes the one 2023 study (extracted in October 2022) to avoid trend confusion.
As to the dataset domains, Table 2 suggests that among the 630 unique “study - dataset domain” pairs, the majority (70.95%, N=447) are in the “product/service review” category, followed by 15.56% (N=98) in “Non-specific”. The third place is shared by two magnitude-smaller categories: “student feedback/ education review” (3.02%, N=19) and “video/movie review” (3.02%, N=19). The numbers of studies with datasets from the Healthcare/medicine (1.43%, N=9) and Politics/policy-reaction (0.79%, N=5) domains were again single-digit. Moreover, nearly half of the unique datasets in the public domains were created by the authors for the first time: 5/9 in Healthcare/medicine, 2/4 in Politics/policy-reaction, and 8/12 in Student feedback/ Education review.
Furthermore, to understand the dataset diversity across samples and domains, we grouped the 1179 unique “study - dataset” pairs by “research-domain, dataset-domain, dataset-name” combinations and zoomed into the 757 entries with ten or more study counts each. As shown in Table 3 and illustrated in Figure 4, among these 757 unique combinations, 95.77% (N=725) are in the “non-specific” research domain, of which 90.48% (N=656) used “product/service review” datasets. Most interestingly, these 757 entries only involve 12 distinct datasets of which 10 were product and service reviews, and 78.20% (N=592) are taken up by the four SemEval datasets from the early pioneer work [31, 32, 33] mentioned in Subsection 1.1: SemEval 2014 Restaurant, SemEval 2014 Laptop (these two alone account for 50.33% of all 757 entries), SemEval 2016 Restaurant, and SemEval 2015 Restaurant. This finding echos [96, 36]: “The SemEval challenge datasets … are the most extensively used corpora for aspect-based sentiment analysis” [36,  p.4]. Meanwhile, the top dataset used under “product/service review” research and dataset domains is the original product review dataset created by the researchers. [36, 97] provides a detailed introduction to the SemEval datasets.
Table 3: Number of Studies Per Each Research (Application) Domain, Dataset Domain, and Dataset Combination for All Datasets Used by Ten or More In-Scope Studies (N=757)
Research Domain Dataset Domain Dataset Count of Studies % of Studies
non-specific product/service review SemEval 2014 Restaurant 200 26.42%
non-specific product/service review SemEval 2014 Laptop 181 23.91%
non-specific product/service review SemEval 2016 Restaurant 110 14.53%
non-specific product/service review SemEval 2015 Restaurant 101 13.34%
non-specific non-specific Twitter (Dong et al., 2014) 53 7.00%
non-specific product/service review Amazon customer review datasets (Hu & Liu, 2004) 25 3.30%
product/service review product/service review Product review (original) 17 2.25%
non-specific non-specific Twitter (original) 16 2.11%
non-specific product/service review SemEval 2015 Laptop 16 2.11%
product/service review product/service review Amazon product review (original) 15 1.98%
non-specific product/service review Yelp Dataset Challenge Reviews 12 1.59%
non-specific product/service review SemEval 2016 Laptop 11 1.45%
Total 757 100.00%
Figure 4: Number of Studies Per Each Research (Application) Domain (left), Dataset Domain (middle), and Dataset (right) Combination, filtered by Datasets Used by 10 or More In-Scope Studies (N=757). The three-way relationship highlights that not only did the majority of the sample studies with “non-specific” research domain use datasets from the ‘product/service review‘ domain, but their datasets were also dominated by only four SemEval datasets on two types of product and service reviews.
It is noteworthy that among the 519 reviewed studies, 20 focused on cross-domain or domain-agnostic ABSA, and 19 of them did not have a specific research application domain. However, while all 20 studies used multiple datasets, 17 solely involved the “product/service review” domain category by using reviews of restaurants and different products, and 14 used at least one SemEval dataset. The only three studies that went beyond the “product/service review” dataset domain added in movie reviews, singer reviews, and generic tweets.
4.2 Results for RQ2
RQ2. What are the most common ABSA problem formulations via subtask combinations, and what proportion of ABSA studies only focus on a specific sub-task?
For RQ2, we examined the 13 recorded subtasks and 807 unique “study - subtask” pairs to identify the most explored ABSA subtasks and subtask combinations across the 519 reviewed studies. As shown in Table 4, 32.37% (N=168) of the studies developed full-ABSA solutions through the combination of AE and ASC, and a similar proportion (30.83%, N=160) focused on ASC alone, usually formulating the research problem as contextualised sentiment analysis with given aspects and the full input text. Only 15.22% (N=79) of the studies solely explored the AE problem. This is consistent with the number of studies by individual subtasks shown in Table 5, where ASC is the most explored subtask, followed by AE and ACD. Moreover, Figure 5 reveals a small but noticeable rise in composite subtasks ASTE (N=1, 5 and 10 in 2017, 2021, 2022) since 2020 and a decline in ASC and AE around the same period.
Table 4: Number of Studies by ABSA Subtask Combinations (N=519)
Subtask Combination Count of Studies % of Studies
AE, ASC 168 32.37%
ASC 160 30.83%
AE 79 15.22%
ACD, ASC 16 3.08%
ASTE 15 2.89%
AE, ACD, ASC 13 2.50%
AE, OE, ASC 9 1.73%
AE, ACD 9 1.73%
AOPE 9 1.73%
ACD 7 1.35%
AE, OE 7 1.35%
AE, OE, ACD 4 0.77%
AE, OE, ACD, ASC 4 0.77%
ASQE 2 0.39%
AOPE, ASC 2 0.39%
OE 1 0.19%
OTE (opinion target extraction), ASC 1 0.19%
Aspect-based embedding, ACD, ASC 1 0.19%
ASC, OE 1 0.19%
Aspect and synthetic sample discrimination 1 0.19%
AE, OE, ASC, AOPE, ASTE 1 0.19%
AOPE, ASC, ACD 1 0.19%
AOPE, ACD 1 0.19%
AE, review-level SA 1 0.19%
ACD, AE, ASC 1 0.19%
AE, ASC, Aspect-based sentence segmentation 1 0.19%
AE, ASC, Aspect-based embedding 1 0.19%
AE, ASC, ACD 1 0.19%
AE, ACD, OE 1 0.19%
Data augmentation 1 0.19%
TOTAL 519 100.00%
Table 5: Number of Studies by Individual ABSA Subtasks (N=807)
Individual Subtask Count of Studies % of Studies
ASC 381 47.21%
AE 300 37.17%
ACD 59 7.31%
OE 28 3.47%
ASTE 16 1.98%
AOPE 14 1.73%
ASQE 2 0.25%
Aspect-based embedding 2 0.25%
Aspect and synthetic sample discrimination 1 0.12%
OTE (opinion target extraction) 1 0.12%
Aspect-based sentence segmentation 1 0.12%
Data augmentation 1 0.12%
Review-level SA 1 0.12%
TOTAL 807 100.00%
Figure 5: Distribution of Unique “Study - ABSA Subtask” Pairs by Publication Year (N=805). This graph excludes the one 2023 study (extracted in October 2022) to avoid trend confusion.
4.3 Results for RQ3
RQ3. What is the trend in the ABSA solution approaches over time? Are linguistic and traditional machine-learning approaches still in use?
To answer RQ3, we examined the 519 in-scope studies along two dimensions, which we call “paradigm” and “approach”. We use “paradigm” to indicate whether a study employed techniques along the supervised-unsupervised dimension and other types, such as reinforcement learning. We classify non-machine-learning approaches under the “unsupervised” paradigm, as our focus is on dataset and resource dependency. By “approach”, we refer to the more specific type of techniques, such as deep learning (DL), traditional machine learning (traditional ML), linguistic rules (“rules” for short), syntactic features and relations (“syntactics” for short), lexicon lists or databases (“lexicon” for short), and ontology or knowledge-driven approaches (“ontology” for short).
Overall, the results suggest that our samples are dominated by fully- (60.89%) and partially-supervised (5.40%) ML methods that are more reliant on annotated datasets and prone to their impact. As to ABSA solution approaches, the sample shows that DL methods have rapidly overtaken traditional ML methods since 2017, particularly with the prevalent RNN family (55.91%) and its combination with the fast-surging attention mechanism (26.52%). Meanwhile, traditional ML and linguistic approaches have remained a small but steady force even in the most recent years. Context engineering through introducing linguistic and knowledge features to DL and traditional ML approaches was very common. More detailed results and richer findings are presented below.
4.3.1 Paradigms
Table 6 lists the number of studies per each of the main paradigms. Among the 519 reviewed studies, 66.28% (N= 344) is taken up by those using somewhat- (i.e. fully-, semi- and weakly-) supervised paradigms that have varied levels of dependency on labelled datasets, where the fully-supervised ones alone account for 60.89% (N=316). Only 19.65% (N=102) of the studies do not require labelled data, which are mostly unsupervised (18.69%, N=97). In addition, hybrid studies are the third largest group (14.07%, N=73).
Table 6: Number of Studies by Paradigms (N=519). Note the unsupervised category includes non-machine learning approaches)
Paradigm Count of Studies % of Studies
Fully-supervised 316 60.89%
Unsupervised 97 18.69%
Hybrid 73 14.07%
Semi-supervised 22 4.24%
Weakly-supervised 6 1.16%
Reinforcement learning 3 0.58%
Self-supervised 2 0.39%
TOTAL 519 100.00%
We further analysed the approaches under each paradigm and focused on three for more details: deep learning (DL), traditional machine learning (ML), and Linguistic and Statistical Approaches. The results are displayed in Tables 7 to  10 and described below.
Table 7: Number of Studies by Deep Learning (DL) Approaches (N=313)
DL Approach Count of Studies % of Studies
RNN, Attention 83 26.52%
Attention 60 19.17%
RNN 31 9.90%
RNN, CNN, Attention 25 7.99%
CNN 20 6.39%
RNN, CNN 19 6.07%
GNN/GCN, Attention 17 5.43%
CNN, Attention 17 5.43%
RNN, GNN/GCN, Attention 17 5.43%
Other (<5% each) 24 7.67%
TOTAL 313 100.00%
Table 8: Number of Studies by Paradigm and Approaches (N=519)
Paradigm Approaches Count of Studies % of Studies
supervised DL 149 47.15%
Syntactics, DL 53 16.77%
DL, traditional ML 23 7.28%
Other (<5% each) 91 28.80%
Total 316 100.00%
semi-supervised DL 7 31.82%
Syntactics, Lexicon, traditional ML 3 13.64%
traditional ML 2 9.09%
Rules, Syntactics, Lexicon 2 9.09%
Other (<5% each) 8 36.36%
Total 22 100.00%
weakly-supervised DL, traditional ML 2 33.33%
traditional ML 1 16.67%
DL 1 16.67%
Syntactics, DL 1 16.67%
Syntactics, Ontology, DL 1 16.67%
Other (<5% each) 0 0.00%
Total 6 100.00%
self-supervised DL 2 100.00%
Other (<5% each) 0 0.00%
Total 2 100.00%
unsupervised Rules, Syntactics, Lexicon 24 24.74%
Rules, Syntactics, Lexicon, Ontology 15 15.46%
Rules, Syntactics 8 8.25%
traditional ML 7 7.22%
Other (<5% each) 43 44.33%
Total 97 100.00%
hybrid Rules, Syntactics, traditional ML 6 8.22%
Rules, Syntactics, Lexicon, traditional ML 6 8.22%
traditional ML 5 6.85%
Syntactics, traditional ML 4 5.48%
Rules, Syntactics, Lexicon 4 5.48%
Rules, Syntactics, Lexicon, Ontology, traditional ML 4 5.48%
Other (<5% each) 44 60.27%
Total 73 100.00%
reinforcement learning DL 1 33.33%
Syntactics, DL 1 33.33%
Other (<5% each) 1 33.33%
Total 3 100.00%
TOTAL 519 100.00%
4.4 Approaches
As shown in Tables 7 and  8, among the 519 reviewed studies, 60.31% (N=313) employed DL approaches, and 30.83% (N=160) are DL-only. The DL-only approach is particularly prominent among fully-supervised (47.15%, N=149) and semi-supervised (31.82%, N=7) studies. Supplementing DL with syntactical features is also the second most popular approach in fully-supervised studies (16.77%, N=53).
  1. 1) DL Approaches
Table 7 suggests that the 313 studies involving DL approaches are dominated by Recurrent Neural Network (RNN)-based solutions (55.91%, N=175), of which nearly half used a combination of RNN and the attention mechanism (26.52%, N=83), followed by attention-only (19.17%, N=60) and RNN-only (9.90%, N=31) models. In addition, convolutional and graph-neural approaches (e.g. Convolutional Neural Networks (CNN), Graph Neural Networks (GNN), Graph Convolutional Networks (GCN)) also play smaller but noticeable roles in DL-based ABSA studies, with the former commonly used as an alternative to the sequence models [20, 98, 99] and the latter for injecting external knowledge and relations to the architecture or better capture local and global context [86, 100, 101].
Figure 6 depicts the trend of the main approaches across the publication years. We excluded the one study pre-published for 2023 to avoid confusing trends. It is clear that DL approaches have risen sharply and taken dominance since 2017, mainly driven by the rapid growth in RNN- and attention-based studies. GNN/GCN-based approaches remain small in number but have noticeable growth since 2020 (N=2, 2, 16, 24 in each of 2019-2022, respectively).
Figure 6: Distribution of Studies Per Method by Publication Year (N=1017 with 519 unique studies). This graph excludes the one 2023 study (extracted in October 2022) to avoid trend confusion.
Table 9: Number of Studies by Traditional Machine Learning Approaches (N=283)
Traditional ML Count of Studies % of Studies
Support Vector Machine (SVM) 57 20.14%
Conditional Random Field (CRF) 41 14.49%
Latent Dirichlet allocation (LDA) 36 12.72%
Naïve Bayes (NB) 32 11.31%
Random Forest (RF) 24 8.48%
Decision Tree (DT) 18 6.36%
Logistic Regression (LR) 15 5.30%
K-Nearest Neighbors (KNN) 12 4.24%
Multinomial Naïve Bayes (MNB) 10 3.53%
K-means 5 1.77%
Boosting 5 1.77%
Other (N<5) 28 9.89%
TOTAL 283 100.00%
  1. 2) Traditional ML Approaches
Interestingly, as shown in Figure 6 traditional ML approaches remain a steady force over the decades despite the rapid rise of DL methods. Tables 8 and  9 provide some insight into this: Among the 519 reviewed studies, while 60.31% employed DL approaches as mentioned in the previous sub-section, over half (54.53%, N=283) also included traditional ML approaches, with the top 3 being Support Vector Machine (SVM; 20.14%, N=57), Conditional Random Field (CRF; 14.49%, N=41), and Latent Dirichlet allocation (LDA; 12.72%, N=36). Table 8 suggests that among the major paradigms, traditional ML were often used in combination with DL approaches for fully-supervised studies (7.28%, N=23), and along with linguistic rules, syntactic features, and/or lexicons and ontology in hybrid studies (27.40%, N=20). Across all paradigms, traditional ML-only approaches are relatively rare (max N=7).
  1. 3) Linguistic and Statistical Approaches
While Table 8 illustrates the prevalence of fusing ML approaches with linguistic and statistical features or modules, there were 67 studies (12.91% out of the total 519) on pure linguistic or statistical approaches. As shown in Tables 10, although small in number, these non-ML approaches have persisted over time. The most popular combination (34.33%, N=23) was rules built on syntactic features (e.g. POS tags and dependency parse trees) and used along lexicon resources (e.g. domain-specific aspect lists, SentiWordNet [^3] https://github.com/aesuli/SentiWordNet, MPQA [^4] https://mpqa.cs.pitt.edu/). A typical example is using POS tags and/or lexicon resources to narrow the scope of aspect or opinion term candidates, applying further rules based on POS tags or dependency relations for AE or OE, and/or using lexicon resources for candidate pruning, categorisation, or sentiment labelling. The second top combination (23.88%, N=16) is the above-mentioned one plus ontology (e.g. domain-specific ontology, ConceptNet [^5] https://conceptnet.io/, WordNet [^6] https://wordnet.princeton.edu/) to bring in external knowledge of concepts and relations. Pure statistical methods were relatively rare (5.97%, N=4), and mainly included frequency-based methods such as N-gram and TF-IDF, and other statistical modelling methods that were not commonly seen in the ML field.
Table 10: Number of Studies by Pure Linguistic or Statistical Approaches and Publication Year (N=67)
Approach 2011-2013 2014-2016 2017-2019 2020-2022 Total Total%
Syntactics, Rules, Lexicon 1 8 8 6 23 34.33%
Syntactics, Rules, Lexicon, Ontology 1 3 9 3 16 23.88%
Syntactics, Rules 2 4 5 11 16.42%
Statistical 2 1 1 4 5.97%
Syntactics, Rules, Lexicon, Statistical 1 1 1 3 4.48%
Syntactics, Rules, Statistical 2 1 3 4.48%
Lexicon 1 1 1.49%
Rules 1 1 1.49%
Rules, Lexicon 1 1 1.49%
Rules, Lexicon, Statistical 1 1 1.49%
Syntactics, Lexicon 1 1 1.49%
Syntactics, Lexicon, Statistical 1 1 1.49%
Syntactics, Statistical 1 1 1.49%
TOTAL 6 16 26 19 67 100.00%
5 Discussion
This review was motivated by the literature gap in capturing trends in ABSA research to answer higher-level questions beyond technical details, and the concern that the domain-dependent nature could predispose ABSA research to systemic hindrance from a combination of resource-reliant approaches and skewed resource domain distribution. By systematically reviewing the 519 in-scope primary studies published between 2008 and 2022, our quantitative analysis results identified trends in ABSA solution approaches, confirmed the above-mentioned concern, and provided detailed insights into the relevant issues. In this section, we examine the primary findings, share ideas for future research, and reflect on the limitations of this review.
5.1 Significant Findings and Trends
5.1.1 The Out-of-sync Research and Dataset Domains
Under RQ1, we examined the distributions of and relationships between our sample’s research (application) domains and dataset domains. The results showed strong skewness in both types of domains and a significant mismatch between them: While the majority (65.32%, N=339) of the 519 studies did not aim for a specific research domain, a greater proportion (70.95%, N=447) used datasets from the “product/service review” domain. A closer inspection of the link between the two domains revealed a clear mismatch: Among the 757 unique “research-domain, dataset-domain, dataset-name” combinations with ten or more studies: 90.48% (N=656) of the studies in the “non-specific” research domain (95.77%, N=725) used datasets from the “product/service review” domain. This suggests that the lack of non-commercial-domain datasets could have forced generic technical studies to use benchmark datasets from a single popular domain. Given ABSA problem’s domain-dependent nature, this could have indirectly hindered the solution development and evaluation across domains.
The results also showed that the other important and prevalent ABSA application domains such as education, medicine/healthcare, and public policy, were clearly under-researched and under-resourced. Among the reviewed samples from these three public-sector domains, about half of their datasets were created for the studies by their authors, indicating a lack of public dataset resources, hence the cost and challenge of developing ABSA research in these areas. As a likely consequence, even the most researched domain among these three had only 12 studies (2.31% out of 519) since 2008. The dataset resource scarcity in these public sector domains deserves more research community attention and support, especially given these domains’ overall low research resources vs. the high cost and domain knowledge required for quality data annotation. In particular, for domains such as “Student feedback/education review” that often face strict data privacy and consent restrictions, it is crucial that the ABSA research community focus on creating ethical and open-access datasets to leverage community resources.
5.1.2 The Dominance and Limitations of the SemEval Datasets
The results under RQ1 also revealed further issues with dataset diversity, even within the dominant “product/service review” domain. Out of the 757 unique “research-domain, dataset-domain, dataset-name” combinations with ten or more studies, 78.20% (N=592) are taken up by the four popular SemEval datasets: The SemEval 2014 Restaurant and Laptop datasets alone account for 50.33% of all 757 entries, and the other two (SemEval 2015 and 2016 Restaurant).
The level of dominance of the SemEval datasets is alerting, not only because of their narrow domain range, but also for the inheritance and impact of the SemEval datasets’ limitations. Several studies (e.g. [36, 96, 102, 94]) suggest that these datasets fail to capture sufficient complexity and granularity of the real-world ABSA scenarios, as they primarily only include single-aspect or multi-aspect-but-same-polarity sentences, and thus mainly reflect sentence-level ABSA tasks and ignored subtasks such as multi-aspect multi-sentiment ABSA. The experiment results from [96, 102, 94] consistently showed that all 35 ABSA models (including those that were state-of-the-art at the time) (9 in [96], 16 in [102], 10 in [94]) that were trained and performed well on the SemEval 2014 ABSA datasets showed various extents of performance drop (by up to 69.73% in [96]) when tested on same-source datasets created for more complex ABSA subtasks and robustness challenges. Given that the SemEval datasets are heavily used as both training data and “benchmark” to measure ABSA solution performance, their limitations and prevalence are likely to form a self-reinforcing loop that confines ABSA research. To break free from this dataset-performance self-reinforcing loop, it is critical that the ABSA research community be aware of this issue, and develop and adopt datasets and practices that are robustness-oriented, such as the automatic data-generation framework and the resulting Aspect Robustness Test Set (ARTS) developed by [96] for probing model robustness in distinguishing target and non-target aspects, and the Multi-Aspect Multi-Sentiment (MAMS) dataset created by [102] to reflect more realistic challenges and complexities in aspect-term sentiment analysis (ATSA) and aspect-category sentiment analysis (ACSA) tasks.
5.1.3 The Reliance on Labelled-Datasets
The domain and dataset issues discussed above would not be as problematic if most ABSA studies employed methods that are dataset-agnostic. However, our results under RQ3 show the opposite. Only 19.65% (N=102, with 97 being unsupervised) of the 519 reviewed studies do not require labelled data, whereas 66.28% (N= 344) are somewhat-supervised, and fully-supervised studies alone account for 60.89% (N=316).
As demonstrated in Section 2.1.2, the domain can directly affect whether a chunk of text is considered an aspect or the relevant sentiment term, and plays a crucial role in contextual inferences such as implicit aspect extraction and multi-aspect multi-sentiment pairing. The domain knowledge reflected via ABSA labelled datasets can further shape the linguistic rules, lexicons, and knowledge graphs for non-machine-learning approaches; and define the underpinning feature space, representations, and acquired relationships and inferences for trained machine-learning models. When applying a solution built on datasets from a domain that is very remote from or much narrower than the intended application domain, it is predictable that the solution performance would be capped at subpar and even fail at more context-heavy tasks [92, 21, 11, 88, 93, 82, 19]. Thus, domain transfer is crucially necessary for balancing the uneven ABSA research and resource distributions across domains. However, our finding that 17 out of the 20 reviewed cross-domain or domain-agnostic ABSA studies solely used datasets from the “product/service review” domain raised questions about these approaches’ generalisability and robustness in other domains, as well as whether such dataset choices became another reinforcement of concentrating research effort and benchmarks within this one dominant domain.
The rapid rise of deep learning (DL) in ABSA research could further add to the challenge of overcoming the negative impact of this domain mismatch and dataset limitations via the non-linear multi-layer dissemination of bias in the representation and learned relations, thus making problem-tracking and solution-targeting difficult. In reality, of the 519 reviewed studies, 60.31% (N=313) employed DL approaches, and nearly half (47.15%, N=149) of the fully-supervised studies and 30.83% (N=160) of all reviewed studies were DL-only.
Moreover, RNN-based solutions dominate the DL approaches (55.91%, N=175), mainly with the RNN-attention combination (26.52%, N=83) and RNN-only (9.90%, N=31) models. RNN and its variants such as Long short-term memory (LSTM) and Gated recurrent units (GRUs) are known for their limitations in capturing long-distance relations due to their sequential nature and the subsequent memory constraints [43, 20]. Although the addition of the attention mechanism enhances the model’s focus on more important features such as aspect terms [43, 20], traditional attention weights calculation struggles with multi-word aspects or multi-aspect sentences [20, 103]. In addition, whilst 16.77% (N=53) of the fully-supervised studies introduced syntactical features to their DL solutions, additional features also increased the input size. According to [104], sequential models, even the state-of-the-art large language models, showed impaired performance as the input grew longer and could not always benefit from additional features.
5.2 Ideas for Future Research
Overall, by adopting a “systematic perspective, i.e., model, data, and training” [94,  p.28] combined with a quantitative approach, we identified high-level trends unveiling the development and direction of ABSA research, and found clear evidence of large-scale issues that affect the majority of the existing ABSA research. The skewed domain distributions of resources and benchmarks could also restrict the choice of new studies. On the other hand, this evidence also highlights areas that need more attention and exploration, such as ABSA solutions and resource development for the less-studied domains (e.g. education and public health), low-resource and/or data-agnostic ABSA, domain adaptation, alternative training schemes such as adversarial (e.g. [94, 105]) and reinforcement learning (e.g. [106, 107]), and more effective feature and knowledge injection.
In addition, our results also revealed emerging trends and new ideas. The relatively recent growth of end-to-end models and composite ABSA subtasks provide opportunities for further exploration and evaluation. The fact that hybrid approaches with non-machine-learning techniques and non-textual features remain steady forces in the field after nearly three decades suggests valuable characteristics that are worth re-examining under the light of new paradigms and techniques. Moreover, this review was conducted before the mass popularity of generative language models, such as ChatGPT [108]. It is worth exploring how ABSA research could ethically leverage the capabilities of pre-trained generative models, potentially for both ABSA solution and resource creation.
Lastly, it is crucial that the community invest in solution robustness, especially for machine-learning approaches [96, 102, 94]. This could mean critical examination of the choice of evaluation metrics, tasks, and benchmarks, and being conscious of their limitations vs. the real-world challenges. The “State-Of-The-Art” (SOTA) performance based on certain benchmark datasets should never become the motivation and holy grail of research, especially in fields like ABSA where the real use cases are often complex and even SOTA models do not generalise far beyond the training datasets. More attention and effort should be paid to analysing the limitations and mistakes of ABSA solutions, and drawing from the ideas of other disciplines and areas to fill the gaps.
5.3 Limitations
We acknowledge the following limitations of this review: First, our sample scope is by no means exhaustive, as it only includes primary studies from four peer-reviewed digital databases and only those published in the English language. Although this can be representative of a core proportion of ABSA research, it does not generalise beyond this without assumptions. Second, no search string is perfect. Our database search syntax and auto-screening keywords represent our best effort in capturing ABSA primary studies, but may have missed some relevant ones, especially with the artificial “total pages <<< 3” and “total keyword (except SA, OM) outside Reference <<< 5” exclusion criteria. Third, we may have missed datasets, paradigms, and approaches that are not clearly described in the primary studies, and our categorisation of them is also subject to the limitations of our knowledge and decisions. Fourth, we did not compare solution performance across studies given the focus of this review and the scale of our sample.
6 Conclusion
ABSA research is riding the wave of the explosion of online digital opinionated text data and the rapid development of NLP resources and ideas. However, its context- and domain-dependent nature and the complexity and inter-relations among its subtasks pose challenges to improving ABSA solutions and applying them to a wider range of domains. In this review, we systematically examined existing ABSA literature in terms of their research application domain, dataset domain, and research methodologies. The results suggest a number of potential systemic issues in the ABSA research literature, including the predominance of the “product/service review” dataset domain among the majority of studies that did not have a specific research application domain, coupled with the prevalence of dataset-reliant methods such as supervised machine learning. We discussed the implication of these issues to ABSA research and applications, as well as their implicit effect in shaping the future of this research field through the mutual reinforcement between resources and methodologies. We suggested areas that need future research attention and proposed ideas for exploration.
Authors’ contributions Y. C. H designed, conducted, and wrote this review. P. D., J. W., and K. T. guided the design of the review methodology and review protocol, and reviewed and provided feedback on the manuscript.
Declarations
Competing Interests The authors have no competing interests as defined by Springer, or other interests that might be perceived to influence the results and/or discussion reported in this paper.
Appendix A Appendix - Full SLR Methodology
This section provides a complete, detailed description of the SLR methodology and procedures.
A.1 Research identification
To obtain the files for review, we conducted database searches between 24-25 October 2022, when we manually queried and exported a total of 4191 research papers’ PDF and BibTeX (or the equivalent) files via the web interfaces of four databases. Table A.1 details the search string, search criteria, and the PDF files exported from each database.
Table A.1: Digital databases and search details used for this Systematic Literature Review (SLR)
Database Search String Search Criteria PDFs exported Type of exported publications
ACM Digital Library ((“aspect based” OR “aspect-based”) AND (sentiment OR extraction OR extract OR mining)) OR “opinion mining” No year filter Search scope = Full article Content Type = Research Article Media Format = PDF Publications = Journals OR Proceedings OR Newsletters 1514 201 articles 1283 conference papers 30 newsletters
IEEE Xplore ((“aspect based” OR “aspect-based”) AND (sentiment OR extraction OR extract OR mining)) OR “opinion mining” Year filter = 2004-2022 (pilot search suggested that 1995-2003 results were irrelevant) Search scope = Full article Publications = Not “Books” 1639 165 articles 1445 conference papers 29 magazine pieces
Science Direct (“aspect based” OR “aspect-based”) AND (sentiment OR extraction OR extract OR mining)) OR “opinion mining” No year filter Search scope = Title, abstract or author-specified keywords Publications = Research Articles only 497 497 articles
Springer Link (“aspect based” OR “aspect-based”) AND (sentiment OR extraction OR extract OR mining)) OR “opinion mining” No year filter English results only Publications = Article, Conference Paper 541 218 articles 323 conference papers
Given the limited search parameters allowed in these digital databases, we adopted a “search broad and filter later” strategy. These database search strings were selected based on pilot trials to capture the ABSA topic name, the relatively prevalent yet unique ABSA subtask term (“extraction”), and the interchangeable use between ABSA and opinion mining; while avoiding generating false positives from the highly active, broader field of SA. The “filter later” step was carried out during the “selection of primary studies” stage introduced in the next section, which aimed at excluding cases where the keywords are only mentioned in the reference list or sparsely mentioned as a side context, and opinion mining studies that were at document or sentence levels.
A.2 Selection of primary studies
After obtaining the 4191 initial search results, we conducted a pilot manual file examination of 100 files to refine the pre-defined inclusion and exclusion criteria. We found that some search results only contained the search keywords in the reference list or Appendix, which was also reported in [104]. In addition, there are a number of papers that only mentioned ABSA-specific keywords in their literature review or introduction sections, and the studies themselves were on coarser-grained sentiment analysis or opinion mining. Lastly, there were instances of very short research reports that provided insufficient details of the primary studies. Informed by these observations, we refined our inclusion and exclusion criteria to those in Table 1 in Section 3. Note that we did not include popularity criteria such as citation numbers so we can better identify novel practices and avoid mainstream method over-dominance introduced by the citation chain [109].
To implement the inclusion and exclusion criteria, we first applied PDF mining to automatically exclude files that meet the exclusion criteria, and then refined the selection with manual screening under the exclusion and inclusion criteria. Both of these processes are detailed below.
The automatic screening consists of a pipeline with two Python packages: Pandas [110] and PyMuPDF [^7] https://pypi.org/project/PyMuPDF/. We first used Pandas to extract into a dataframe (i.e. table) all exported papers’ file locations and key BibTex or equivalent information including title, year, page number, DOI, and ISBN. Next, we used PyMuPDF to iterate through each PDF file and add to the dataframe multiple data fields: whether the file was successfully decoded [^8] https://pdfminersix.readthedocs.io/en/latest/faq.html for text extraction (if marked unsuccessful, the file was marked for manual screening), the occurrence count of each Regex keyword pattern listed below, and whether each keyword occurs after the section headings that fit into Regex patterns that represent variations of “references” and “bibliography” (referred to as “non-target sections” below). We then marked the files for exclusion by evaluating the eight criteria listed under “Auto-excluded” in Table A.2 against the information recorded in the dataframe. Each of the auto-exclusion results from Steps 1-4 and 7 in Table A.2 were manually checked, and those under Steps 5, 6, and 8 were spot-checked. These steps excluded 3277 out of the 4194 exported files.
Below are the regex patterns used for automatic keyword extraction and occurrence calculation:
[…]
For the 914 files filtered through the auto-exclusion process, we manually screened them individually according to the inclusion and exclusion criteria. As shown in the second half of Table A.2, this final screening step refined the review scope to 519 papers.
[b]
Table A.2: The automatic and manual screening processes
Inclusion/Exclusion Types File Count
Total Extracted Files 4191
Auto-excluded 3277
      Step_1 - Duplicate DOI across databases 19
      Step_2 - Duplicate Title across databases 11
      Step_3 - Survey/Not primary study papers 149
      Step_4 - Total keyword outside reference match = 0 26
      Step_5 - keyword matched only to SA and/or OM outside reference 1 2235
[…]
      Step_7 - Total pages <3 7
      Step_8 - Total keyword (except SA, OM) outside Reference < 5 1,2 587
Manually Excluded 395
      Type - Article withdrawn 1
      Type - Not published in English 1
      Type - Dataset unclear 1
      Type - Low quality, unclear definition of ABSA 1
      Type - No original method 2
      Type - Duplicate of another included paper 3
      Type - Sentence level SA 4
      Type - Not text-data focused 6
      Type - Not primary study 15
      Type - Lack of details on ABSA tasks 31
      Type - Review (Not primary study) 43
      Type - No ABSA task experiment details/results 49
      Type - Not ABSA focused; No ABSA task experiment details/results 238
Final Included for Review 519
  • 1
[…]
  • 2
    The occurrence threshold 5 was chosen based on pilot file examination, which suggested that files with target keyword occurrence below this threshold tended to be non-ABSA-focused.
A.3 Data extraction and synthesis
In the final step of the SLR, we manually reviewed each of the 519 in-scope publications and recorded information according to a pre-designed data extraction form. The key information recorded includes each study’s research focus, research application domain (“research domain” below), ABSA subtasks involved, name or description of all the datasets directly used, model name (for machine-learning solutions), architecture, whether a certain approach or paradigm is present in the study (e.g. supervised learning, deep learning, end-to-end framework, ontology, rule-based, syntactic-components), and the specific approach used (e.g. attention mechanism, Naïve Bayes classifier) under the deep learning and traditional machine learning categories.
After the data extraction, we performed data cleaning to identify and fix recording errors and inconsistencies, such as data entry typos and naming variations of the same dataset across studies. Then we created two mappings for the research and dataset domains described below.
For each reviewed study, its research domain was defaulted to “non-specific” unless the study mentioned a specific application domain or use case as its motivation, in which case that domain description was recorded instead.
The dataset domain was recorded and processed at the individual dataset level, as many reviewed studies used multiple datasets. We standardised the recorded dataset names, checked and verified the recorded dataset domain descriptions provided by the authors or the source web-pages, and then manually categorised each domain description into a domain category. For published/well-known datasets, we unified the recorded naming variations and checked the original datasets or their descriptions to verify the domain descriptions. For datasets created (e.g. web-crawled) by the authors of the reviewed studies, we named them following the “[source] [domain] (original)” format, e.g. “Yelp restaurant review (original) ”, or “Twitter (original)” if there was no distinct domain, and did not differentiate among the same-name variations. In all of the above cases, if a dataset was not created with a specific domain filter (e.g. general Twitter tweets), then it was classified as “non-specific”.
The recorded research and dataset domain descriptions were then manually grouped into 19 common domain categories. We tried to maintain consistency between the research and dataset domain categories. The following are two examples of possible mapping outcomes:
  1. 1.
    A study on a full ABSA solution without mentioning a specific application domain and using Yelp restaurant review and Amazon product review datasets would be assigned a research domain of “non-specific” and a dataset domain of “product/service review”.
  2. 2.
    A study mentioning “helping companies improve product design based on customer reviews” as the motivation would have a research domain of “product/service review”, and if they used a product review dataset and Twitter tweets crawled without filtering, the dataset domains would be “product/service review” and “non-specific”.
After applying the above-mentioned standardisation and mappings, we analysed the synthesised data quantitatively using the Pandas [110] library to obtain an overview of the reviewed studies and explore the answers to our RQs.
Appendix B Appendix - Additional Results
Figure B.1: Number of Included Studies by Publication Year and Type (N=519)
Note: Although our original search scope included journal articles, conference papers, newsletters, and magazine articles, the final 519 in-scope studies consist of only journal articles and conference papers. Conference papers noticeably outnumbered journal articles in all years until 2022, with the gap closing since 2016. We think this trend could be due to multiple factors, such as the fact that our search was conducted in late October 2022 when some conference publications were still not available; the publication lag for journal articles due to a longer processing period; and potentially a change in publication channels that is outside the scope of this review.